# A Miyaoka-Yau inequality for hyperplane arrangements in $\mathbb{CP}^{n}$

### v1

@@ 1 Introduction
Let $\mathcal{H}$ be an arrangement of complex hyperplanes $H\subset\mathbb{CP}^{n}$. Let $N=|\mathcal{H}|$ be the number of hyperplanes, which we assume to be finite, and fix a labelling, say =$\mathcal{H}=\{H_{1},\ldots,H_{N}\}$. We assume that $H_{i}\neq H_{j}$ for $i\neq j$. In this paper, we introduce a quadratic form $Q:\mathbb{R}^{N}\to\mathbb{R}$ associated to the arrangement $\mathcal{H}$ and show that $Q\leq 0$ on a certain convex polyhedral cone $C\subset\mathbb{R}_{\geq 0}^{N}$, that we call the semistable cone of $\mathcal{H}$.
@@
$Q_{ij}=\begin{cases}-(n+1)\sigma_{i}+2n&\text{ if }i=j\,,\\ -2&\text{ if }i\neq j\text{ and }L=H_{i}\cap H_{j}\text{ is reducible},\\ \,\,n-1&\text{ if }i\neq j\text{ and }L=H_{i}\cap H_{j}\text{ is irreducible}.\end{cases}$
@@
The semistable and stable cones. A basis $\mathcal{B}$ of $\mathcal{H}$ is a subset $\mathcal{B}\subset\mathcal{H}$ of $n+1$ linearly independent hyperplanes. The indicator function $\mathbf{e}_{\mathcal{B}}$ of a basis $\mathcal{B}$ is the vector in $\mathbb{R}^{N}$ whose $i$-th component is $1$ if $H_{i}\in\mathcal{B}$ and 00 otherwise. The semistable cone $C$ is the conical hull of the vectors $\mathbf{e}_{\mathcal{B}}$ with $\mathcal{B}$ ranging over all bases of $\mathcal{H}$. Put differently, if $P$ is the convex hull of the vectors $\mathbf{e}_{\mathcal{B}}$ , then $C=\mathbb{R}_{\geq 0}\cdot P$ is the cone over $P$. The convex set $P$ is called the matroid polytope of $\mathcal{H}$.
@@
Conjecture: suppose that $\mathbf{a}\in C^{\circ}$ is such that $Q(\mathbf{a})=0$ and $10<a_{i}<$ 1 for all $i$. Then there is a Kähler metric on $\mathbb{CP}^{n}$ of constant holomorphic sectional curvature with cone angles $2\pi\alpha_{i}$ along the hyperplanes $H_{i}$ , where $\alpha_{i}=1-a_{i}$. (For a more precise version, in the case of zero curvature, see Conjecture 1.3.)
@@ 1.1 Main result: klt and CY arrangements
[…]
@@
Note that the klt condition (1.1) implies that $10<a_{L}<$ 1.
@@
We calculate $\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ in terms of the intersection poset of the arrangement $\mathcal{H}$. Let $h=c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ be the generator of $H^{2}(\mathbb{CP}^{n},\mathbb{Z})$, and let $\gamma_{L}\in H^{2}(X,\mathbb{Z})$with $L\in\mathcal{L}_{\textup{irr}}$- be the Poincaré dual of the divisor $D_{L}$. We derive the formula
@@
The Picard group of $X$ and the polarization. The Picard group of $X$ is the free abelian group generated by $[\pi^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1)]$ together with the classes $[D_{L}]$ of the exceptional divisors $D_{L}$ of the birational morphism $\pi:X\to\mathbb{CP}^{n}$. The exceptional divisors $D_{L}$ correspond to non-empty irreducible subspaces $L\subset\mathbb{CP}^{n}$ of codimension $\geq 2$, we denote the set of all such subspaces by $\mathcal{L}_{\textup{irr}}^{\circ}:=\mathcal{L}_{\textup{irr}}\setminus\mathcal{H}$.
@@
The above sum is over all hyperplanes $H\in\mathcal{H}$ that are transverse to $\mathcal{V}$. The term $O(k^{n-2})$ denotes a polynomial in $k$ of degree at most $n-2$ whose coefficients are uniformly bounded in absolute value by a number that depends only on the arrangement $\mathcal{H}$ and the polarization coefficients $b_{L}$.
@@
The cup product $p(k)=c_{1}(P_{k})^{n-2}\cdot\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ defines a polynomial of degree $n-2$ in $k$ that we write as
@@
The leading order coefficient of $p(k)$ is given by $C_{n-2}=(\pi^{*}h)^{n-2}\cdot\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ More geometrically, the coefficient $C_{n-2}$ is the parabolic second Chern character of the restriction of $\mathcal{E}_{*}$ to a generic $2$-plane.
@@ 2.1 Basic definitions
Intersection poset. The intersection poset $\overline{\mathcal{L}}$ of $\mathcal{H}$ is the set of all subspaces $L\subset\mathbb{CP}^{n}$ obtained as intersection of hyperplanes in $\mathcal{H}$. We equip $\overline{\mathcal{L}}$ with the partial order given by reverse inclusion. The intersection poset $\overline{\mathcal{L}}$ has a unique minimal element $L=\mathbb{CP}^{n}$ corresponding to the intersection over the empty subset of $\mathcal{H}$ and a unique maximal element $L=T$ given by the centre of the arrangement. We will mainly work with the subposet $\mathcal{L}\subset\overline{\mathcal{L}}$ consisting of all non-empty and proper intersection subspaces. In other words, we exclude $\mathbb{CP}^{n}$ and $\emptyset$ (if $\mathcal{H}$ is essential) from $\overline{\mathcal{L}}$ and let =¯$\mathcal{L}=\overline{\mathcal{L}}\,\setminus\,\{\emptyset,\,\mathbb{CP}^{n}\}$.
@@
$\mathcal{H}_{L}=\{H\in\mathcal{H}\,\text{ such that }\,L\subset H\}.$
@@
$\mathcal{H}^{L}=\{H\cap L\textup{ with }H\in\mathcal{H}\setminus\mathcal{H}_{L}\}.$
@@ 2.2 Irreducible arrangements and subspaces
where $I_{1}$ and $I_{2}$ are subsets of $[n+1]=\{1,\ldots,n+1\}$. The defining linear equation for the hyperplanes in $\mathcal{H}_{1}$ depend only on the variables $x_{i}$ for $i\in I_{1}$ and similarly for $\mathcal{H}_{2}$. We will show that $I_{1}\cap I_{2}=\emptyset$. The subspace $T_{1}$ is the span in projective space of the vectors $\partial_{x_{i}}$ for $i$ in the complement index set $I^{c}_{1}=[n+1]\setminus I_{1}$ and similarly for $T_{2}$. The fact that $T_{1}+T_{2}=\mathbb{CP}^{n}$ implies that $I_{1}^{c}\cup I_{2}^{c}=[n+1]$. Taking complements we get that $I_{1}\cap I_{2}=\emptyset$. The converse is similar. ∎
@@
The product arrangement $\mathcal{H}_{1}\times\mathcal{H}_{2}\subset\mathbb{CP}^{n}$ is obtained by taking the joins $H_{1}+P_{2}$ for $H_{1}\in\mathcal{H}_{1}$ together with $P_{1}+H_{2}$ for $H_{2}\in\mathcal{H}_{2}$
@@ 2.3 Weighted arrangements
Since for $L\in\mathcal{L}$ the multiplicity $m_{L}$ is always greater or equal than $\operatorname{codim}L$, we must have $|\mathcal{H}|>n+1$. Consider the CY weighted arrangement $(\mathcal{H},\mathbf{a})$ where all weights $a_{H}$ are equal to $(n+1)/|\mathcal{H}|$. As $|\mathcal{H}|>n+1$, we have $10<a_{H}<$ 1. Equation (2.8) guarantees that $(\mathcal{H},\mathbf{a})$ is klt. The result follows from Lemma 2.27 . ∎
@@ 3.1 The minimal De Concini-Procesi model
[…]
@@
with components the linear projections $\operatorname{pr}_{L}$
@@
It is shown in [DCP95, §$\S 1.6$] that $D_{L}$ is the closure of $\pi^{-1}(L^{\circ})$ where $L^{\circ}$ is the complement of the induced arrangement $\mathcal{H}^{L}$. Moreover, [DCP95, §$\S 4.3$] shows that $D_{L}$ is biholomorphic to the product
@@ 3.1.1 Construction of $X$ as an iterated blowup
[…]
@@
[…]
@@
[…]
@@ 3.1.3 Example: tetrahedron in $\mathbb{CP}^{3}$
Let $p_{1},\ldots,p_{5}\in\mathbb{CP}^{3}$ be five points in general linear position, i.e., no $3$ points lie on a line and no $4$ points lie on a plane. We can represent these five points as the four vertices of a tetrahedron in $\mathbb{CP}^{3}$ together with its barycentre. Each triplet of points determines a plane, giving rise to a collection of $10$ planes $\mathcal{H}\subset\mathbb{CP}^{3}$see Figure 1 .
[…]
@@
[…]
@@
More generally, the braid or $A_{n+1}$-arrangement in $\mathbb{CP}^{n}$ is the collection of $\binom{n+2}{2}$ hyperplanes spanned by $n$ out of $n+2$ points in general linear position. The arrangement complement $(\mathbb{CP}^{n})^{\circ}$ is naturally identified with the configuration space $\mathcal{M}_{0,n+3}$ of $n+3$ marked points in the Riemann sphere modulo Möbius transformations. It is well known that the minimal De Concini-Procesi model (3.2) agrees with the Deligne-Mumford-Knudsen compactification ¯$\overline{\mathcal{M}}_{0,n+3}$ , see [Kap93, Theorem 4.3.3].
@@ 3.2 Picard group and cohomology of $X$
(ii) Let $s$ be a section of $\mathcal{O}_{X}(D)$ with $D=s^{-1}(0)$. Since $Y\not\subset D$, the zero set of the pullback section $\pi^{*}s$ is equal to $\widetilde{D}$, providing a trivialization of $\pi^{*}(\mathcal{O}_{X}(D))\otimes\mathcal{O}_{\widetilde{X}}(-\widetilde{D})$. Taking isomorphism classes gives (3.5). ∎
@@
The Picard group of the resolution $X$ is the free abelian group generated by the classes $[D_{L}]$ for $L\in\mathcal{L}_{\textup{irr}}^{\circ}$ together with the class of $\pi^{*}\mathcal{O}_{\mathbb{P}^{n}}(1)$.
@@
[…]
@@
Claim. The Picard group of $X_{i}$ is a free abelian group of rank $i+1$ generated by the $\pi_{i}$-exceptional divisors $[D^{i}_{j}]$ for $1\leq j\leq i$ together with the class of $\pi_{i}^{*}\left(\mathcal{O}_{\mathbb{P}^{n}}(1)\right)$.
We prove the claim by induction on $i$. The case $i=0$ follows from the fact that $\operatorname{Pic}(\mathbb{CP}^{n})$ is generated by $\mathcal{O}_{\mathbb{P}^{n}}(1)$. Suppose that $1\leq i\leq k$ and assume that the Picard group of $X_{i-1}$ is the free abelian group generated by the $\pi_{i-1}$-exceptional divisors $[D^{i-1}_{j}]$ for $11\leq j\leq i-$ 1 together with the class of $\pi_{i-1}^{*}\left(\mathcal{O}_{\mathbb{P}^{n}}(1)\right)$. By Lemma 3.20 (i) the Picard group of $X_{i}$ is the free abelian group generated by $\sigma_{i}^{*}(\pi_{i-1}^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1)))=\pi_{i}^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ together with $\sigma_{i}^{*}([D^{i-1}_{j}])$ for $11\leq j\leq i-$ 1 and $[\sigma_{i}^{-1}(\widetilde{L}_{i})]=[D^{i}_{i}]$. Furthermore, if $\widetilde{L}_{i}\subset D^{i-1}_{j}$ for some $11\leq j\leq i-$ 1 then
@@
which contradicts the inclusion preserving property of the labels. Thus, if $11\leq j\leq i-$ 1 then $\widetilde{L}_{i}\not\subset D^{i-1}_{j}$ and by Lemma 3.20 (ii) $\sigma_{i}^{*}([D^{i-1}_{j}])=[\widetilde{D}^{i-1}_{j}]=[D^{i}_{j}]$. This finishes the proof of the claim.
@@
$[D_{H}]=[\pi^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1))]-\sum_{L|L\subsetneq H}[D_{L}]$ (3.6)
@@
where $\gamma_{V}\in H^{2r}(X,\mathbb{Z})$ and $\gamma_{\widetilde{V}}\in H^{2r}(\widetilde{X},\mathbb{Z})$ are the Poincaré duals of $V$ and $\widetilde{V}$
@@
[…]
@@
[…]

### v2

@@ 1 Introduction
Let $\mathcal{H}$ be an arrangement of complex hyperplanes $H\subset\mathbb{CP}^{n}$. Let $N=|\mathcal{H}|$ be the number of hyperplanes, which we assume to be finite, and fix a labelling, say $\mathcal{H}=\{H_{1},\ldots,H_{N}\}$. We assume that $H_{i}\neq H_{j}$ for $i\neq j$. In this paper, we introduce a quadratic form $Q:\mathbb{R}^{N}\to\mathbb{R}$ associated to the arrangement $\mathcal{H}$ and show that $Q\leq 0$ on a certain convex polyhedral cone $C\subset\mathbb{R}_{\geq 0}^{N}$, that we call the semistable cone of $\mathcal{H}$.
@@
[…]
@@
The semistable and stable cones. A basis $\mathcal{B}$ of $\mathcal{H}$ is a subset $\mathcal{B}\subset\mathcal{H}$ of $n+1$ linearly independent hyperplanes. The indicator function $\mathbf{e}_{\mathcal{B}}$ of a basis $\mathcal{B}$ is the vector in $\mathbb{R}^{N}$ whose $i$-th component is $1$ if $H_{i}\in\mathcal{B}$ and 0 otherwise. The semistable cone $C$ is the conical hull of the vectors $\mathbf{e}_{\mathcal{B}}$ with $\mathcal{B}$ ranging over all bases of $\mathcal{H}$. Put differently, if $P$ is the convex hull of the vectors $\mathbf{e}_{\mathcal{B}}$ , then $C=\mathbb{R}_{\geq 0}\cdot P$ is the cone over $P$. The convex set $P$ is called the matroid polytope of $\mathcal{H}$.
@@
Conjecture: suppose that $\mathbf{a}\in C^{\circ}$ is such that $Q(\mathbf{a})=0$ and $0<a_{i}<1$ for all $i$. Then there is a Kähler metric on $\mathbb{CP}^{n}$ of constant holomorphic sectional curvature with cone angles $2\pi\alpha_{i}$ along the hyperplanes $H_{i}$ , where $\alpha_{i}=1-a_{i}$. (For a more precise version, in the case of zero curvature, see Conjecture 1.3.)
@@ 1.1 Main result: klt and CY arrangements
The parabolic bundle. Let $X\xrightarrow{\pi}\mathbb{CP}^{n}$ be the resolution of $\mathcal{H}$ and let $\mathcal{E}=\pi^{*}(T\mathbb{CP}^{n})$ be the pullback tangent bundle. We define a natural parabolic bundle (Definition 4.1) $\mathcal{E}_{*}$ on $(X,D)$ where $D=\pi^{-1}(\mathcal{H})$.
@@
Note that the klt condition (1.1) implies that $0<a_{L}<1$.
@@
We calculate $\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ in terms of the intersection poset of the arrangement $\mathcal{H}$. Let $h=c_{1}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ be the generator of $H^{2}(\mathbb{CP}^{n},\mathbb{Z})$, and let $\gamma_{L}\in H^{2}(X,\mathbb{Z})$ -with $L\in\mathcal{L}_{\textup{irr}}$- be the Poincaré dual of the divisor $D_{L}$. We derive the formula
@@
The Picard group of $X$ and the polarization. The Picard group of $X$ is the free abelian group generated by $\pi^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1)]$ together with the classes $D_{L}]$ of the exceptional divisors $D_{L}$ of the birational morphism $\pi:X\to\mathbb{CP}^{n}$. The exceptional divisors $D_{L}$ correspond to non-empty irreducible subspaces $L\subset\mathbb{CP}^{n}$ of codimension $\geq 2$, we denote the set of all such subspaces by $\mathcal{L}_{\textup{irr}}^{\circ}:=\mathcal{L}_{\textup{irr}}\setminus\mathcal{H}$.
@@
The above sum is over all hyperplanes $H\in\mathcal{H}$ that are transverse to $\mathcal{V}$. The term $O(k^{n-2})$ denotes a polynomial in $k$ of degree at most n2n-2 whose coefficients are uniformly bounded in absolute value by a number that depends only on the arrangement $\mathcal{H}$ and the polarization coefficients $b_{L}$.
@@
The cup product $p(k)=c_{1}(P_{k})^{n-2}\cdot\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ defines a polynomial of degree n2n-2 in $k$ that we write as
@@
The leading order coefficient of $p(k)$ is given by $C_{n-2}=(\pi^{*}h)^{n-2}\cdot\operatorname{par-ch}_{2}(\mathcal{E}_{*})$ . More geometrically, the coefficient $C_{n-2}$ is the parabolic second Chern character of the restriction of $\mathcal{E}_{*}$ to a generic $2$-plane.
@@ 2.1 Basic definitions
Intersection poset. The intersection poset $\overline{\mathcal{L}}$ of $\mathcal{H}$ is the set of all subspaces $L\subset\mathbb{CP}^{n}$ obtained as intersection of hyperplanes in $\mathcal{H}$. We equip $\overline{\mathcal{L}}$ with the partial order given by reverse inclusion. The intersection poset $\overline{\mathcal{L}}$ has a unique minimal element $L=\mathbb{CP}^{n}$ corresponding to the intersection over the empty subset of $\mathcal{H}$ and a unique maximal element $L=T$ given by the centre of the arrangement. We will mainly work with the subposet $\mathcal{L}\subset\overline{\mathcal{L}}$ consisting of all non-empty and proper intersection subspaces. In other words, we exclude $\mathbb{CP}^{n}$ and $\emptyset$ (if $\mathcal{H}$ is essential) from $\overline{\mathcal{L}}$ and let $\mathcal{L}=\overline{\mathcal{L}}\,\setminus\,\{\emptyset,\,\mathbb{CP}^{n}\}$.
@@
[…]
@@
[…]
@@ 2.2 Irreducible arrangements and subspaces
where $I_{1}$ and $I_{2}$ are subsets of $n+1]=\{1,\ldots,n+1\}$. The defining linear equation for the hyperplanes in $\mathcal{H}_{1}$ depend only on the variables $x_{i}$ for $i\in I_{1}$ and similarly for $\mathcal{H}_{2}$. We will show that $I_{1}\cap I_{2}=\emptyset$. The subspace $T_{1}$ is the span in projective space of the vectors $\partial_{x_{i}}$ for $i$ in the complement index set $I^{c}_{1}=[n+1]\setminus I_{1}$ and similarly for $T_{2}$. The fact that $T_{1}+T_{2}=\mathbb{CP}^{n}$ implies that $I_{1}^{c}\cup I_{2}^{c}=[n+1]$. Taking complements we get that $I_{1}\cap I_{2}=\emptyset$. The converse is similar. ∎
@@
The product arrangement $\mathcal{H}_{1}\times\mathcal{H}_{2}\subset\mathbb{CP}^{n}$ is obtained by taking the joins $H_{1}+P_{2}$ for $H_{1}\in\mathcal{H}_{1}$ together with $P_{1}+H_{2}$ for $H_{2}\in\mathcal{H}_{2}$ .
@@ 2.3 Weighted arrangements
Since for $L\in\mathcal{L}$ the multiplicity $m_{L}$ is always greater or equal than $\operatorname{codim}L$, we must have $|\mathcal{H}|>n+1$. Consider the CY weighted arrangement $(\mathcal{H},\mathbf{a})$ where all weights $a_{H}$ are equal to $(n+1)/|\mathcal{H}|$. As $|\mathcal{H}|>n+1$, we have $0<a_{H}<1$. Equation (2.8) guarantees that $(\mathcal{H},\mathbf{a})$ is klt. The result follows from Lemma 2.27 . ∎
@@ 3.1 The minimal De Concini-Procesi model
sends a point $p\in\mathbb{CP}^{n}\setminus L$ to the intersection of $L+p$ with $P_{L}$. The blowup of $\mathbb{CP}^{n}$ along $L$ is the map $B\xrightarrow{\sigma}\mathbb{CP}^{n}$, where $B\subset\mathbb{CP}^{n}\times P_{L}$ is the closure of the graph of $\operatorname{pr}_{L}$ and $\sigma$ is the restriction to $B$ of the projection to the first factor.
@@
with components the linear projections $\operatorname{pr}_{L}$ .
@@
It is shown in [DCP95, $\S 1.6$] that $D_{L}$ is the closure of $\pi^{-1}(L^{\circ})$ where $L^{\circ}$ is the complement of the induced arrangement $\mathcal{H}^{L}$. Moreover, [DCP95, $\S 4.3$] shows that $D_{L}$ is biholomorphic to the product
@@ 3.1.1 Construction of $X$ as an iterated blowup
If $N$ is a complex manifold and $S\subset N$ is a complex submanifold, then the blowup of $N$ along $S$ is a complex manifold $M$ together with a proper holomorphic map $M\xrightarrow{\sigma}N$ such that $\sigma:M\setminus E\cong N\setminus S$ is a biholomorphism and the restriction of $\sigma$ to $E$ is equivalent to the bundle projection $\mathbb{P}(N_{S})\to S$, where $\mathbb{P}(N_{S})$ is the projectivization of the normal bundle $N_{S}$ of $S\subset N$.
@@
Proposition 3.7 guarantees that the end result $X\xrightarrow{\pi}\mathbb{CP}^{n}$ is independent of the labelling, as long as it is compatible with inclusion relations. For example, we can order the elements of $\mathcal{L}_{\textup{irr}}^{\circ}$ by increasing dimension $\mathcal{L}_{\textup{irr}}^{0},\mathcal{L}_{\textup{irr}}^{1},\ldots,\mathcal{L}_{\textup{irr}}^{n-2}$, where the members of $\mathcal{L}_{\textup{irr}}^{i}$ are taken in any order, c.f. [OT07, Theorem 4.2.4].
@@
where $\operatorname{Bl}_{0}\mathbb{C}^{2}\xrightarrow{\sigma}\mathbb{C}^{2}$ is the blowup of $\mathbb{C}^{2}$ at the origin.
@@ 3.1.3 Example: tetrahedron in $\mathbb{CP}^{3}$
Let $p_{1},\ldots,p_{5}\in\mathbb{CP}^{3}$ be five points in general linear position, i.e., no $3$ points lie on a line and no $4$ points lie on a plane. We can represent these five points as the four vertices of a tetrahedron in $\mathbb{CP}^{3}$ together with its barycentre. Each triplet of points determines a plane, giving rise to a collection of $10$ planes $\mathcal{H}\subset\mathbb{CP}^{3}$ -see Figure 1 .
$p_{1}$$p_{3}$$p_{4}$$p_{2}$$p_{5}$$H_{123}$$L_{12}$
@@
$\mathcal{L}_{\textup{irr}}^{2}=\{H_{ijk}=\overline{p_{i}p_{j}p_{k}}\},\hskip 5.69054pt\mathcal{L}_{\textup{irr}}^{1}=\{L_{ij}=\overline{p_{i}p_{j}}\},\hskip 5.69054pt\mathcal{L}_{\textup{irr}}^{0}=\{p_{1},\ldots,p_{5}\}.$
@@
More generally, the braid or $A_{n+1}$-arrangement in $\mathbb{CP}^{n}$ is the collection of $\binom{n+2}{2}$ hyperplanes spanned by $n$ out of $n+2$ points in general linear position. The arrangement complement $(\mathbb{CP}^{n})^{\circ}$ is naturally identified with the configuration space $\mathcal{M}_{0,n+3}$ of $n+3$ marked points in the Riemann sphere modulo Möbius transformations. It is well known that the minimal De Concini-Procesi model (3.2) agrees with the Deligne-Mumford-Knudsen compactification $\overline{\mathcal{M}}_{0,n+3}$ , see [Kap93, Theorem 4.3.3].
@@ 3.2 Picard group and cohomology of $X$
In the rest of the section we let $\mathcal{H}\subset\mathbb{CP}^{n}$ be an arrangement and let $X\xrightarrow{\pi}\mathbb{CP}^{n}$ be the resolution of $\mathcal{H}$ as in Definition 3.1 . Recall that we write $\mathcal{L}_{\textup{irr}}^{\circ}=\mathcal{L}_{\textup{irr}}\setminus\mathcal{H}$.
@@
The Picard group of the resolution $X$ is the free abelian group generated by the classes $D_{L}]$ for $L\in\mathcal{L}_{\textup{irr}}^{\circ}$ together with the class of $\pi^{*}\mathcal{O}_{\mathbb{P}^{n}}(1)$.
@@
We begin by recalling the construction of $X$ by a sequence of blowups as in Proposition 3.7 . The upshot is that $X\xrightarrow{\pi}\mathbb{CP}^{n}$ is equal to $X_{k}\xrightarrow{\pi_{k}}\mathbb{CP}^{n}$ where $X_{i}\xrightarrow{\pi_{i}}\mathbb{CP}^{n}$ are constructed inductively as follows.
@@
Claim. The Picard group of $X_{i}$ is a free abelian group of rank $i+1$ generated by the $\pi_{i}$-exceptional divisors $D^{i}_{j}]$ for $1\leq j\leq i$ together with the class of $\pi_{i}^{*}\left(\mathcal{O}_{\mathbb{P}^{n}}(1)\right)$.
We prove the claim by induction on $i$. The case $i=0$ follows from the fact that $\operatorname{Pic}(\mathbb{CP}^{n})$ is generated by $\mathcal{O}_{\mathbb{P}^{n}}(1)$. Suppose that $1\leq i\leq k$ and assume that the Picard group of $X_{i-1}$ is the free abelian group generated by the $\pi_{i-1}$-exceptional divisors $D^{i-1}_{j}]$ for $1\leq j\leq i-1$ together with the class of $\pi_{i-1}^{*}\left(\mathcal{O}_{\mathbb{P}^{n}}(1)\right)$. By Lemma 3.20 (i) the Picard group of $X_{i}$ is the free abelian group generated by $\sigma_{i}^{*}(\pi_{i-1}^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1)))=\pi_{i}^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1))$ together with $\sigma_{i}^{*}([D^{i-1}_{j}])$ for $1\leq j\leq i-1$ and $\sigma_{i}^{-1}(\widetilde{L}_{i})]=[D^{i}_{i}]$. Furthermore, if $\widetilde{L}_{i}\subset D^{i-1}_{j}$ for some $1\leq j\leq i-1$ then
@@
which contradicts the inclusion preserving property of the labels. Thus, if $1\leq j\leq i-1$ then $\widetilde{L}_{i}\not\subset D^{i-1}_{j}$ and by Lemma 3.20 (ii) $\sigma_{i}^{*}([D^{i-1}_{j}])=[\widetilde{D}^{i-1}_{j}]=[D^{i}_{j}]$. This finishes the proof of the claim.
@@
$D_{H}]=[\pi^{*}(\mathcal{O}_{\mathbb{P}^{n}}(1))]-\sum_{L|L\subsetneq H}[D_{L}]$ (3.6)
@@
[…]
@@
In a neighbourhood of $p$, we can identify the resolution $X\xrightarrow{\pi}\mathbb{CP}^{n}$ with the blowup of $\mathbb{CP}^{n}$ along $L$ so that:
@@
  • the class $(\pi^{*}h)^{n-2}\cdot\gamma_{L}$ is the Poincaré dual of a curve $C=\pi^{-1}(p)$ which is a fibre of the bundle projection $\mathbb{P}(N_{L})\xrightarrow{\pi}L$.